OpenStudy (anonymous):

A particle moves along a straight line with equation of motion s = t^2 + 3t + 1, determine the velocity at t=10 seconds.

OpenStudy (cwrw238):

velocity at time t = ds/dt find ds/dt and plug in t=10

OpenStudy (anonymous):

can you demonstrate using difference quotient?

OpenStudy (cwrw238):

if s = a t^n ds/dt = an t^(n-1)

OpenStudy (anonymous):

yes, what cwrw posted is just the power rule.

OpenStudy (anonymous):

Can you find the derivative? Just apply the power rule to each of the terms. To, \(\large\color{black}{ \tt t^2 }\) , \(\large\color{black}{ \tt 3t }\) and derivative of \(\large\color{black}{ \tt 1 }\), just like derivative of any other constant, is zero.

OpenStudy (anonymous):

@cleetus779 ?

OpenStudy (anonymous):

Anything you don't understand? please reply don't be scilent.

OpenStudy (anonymous):

Simply speaking, you have \(\large\color{black}{ \tt s(t) = t^2 + 3t + 1 }\) And you need, \(\large\color{black}{ \tt s'(10) }\)

OpenStudy (cwrw238):